OpenStudy (anonymous):

Help! Will Medal Please... If enough oxygen is presented for 132g of propane (C3H8) to burn completely in the given reaction what mass of water is produced? Give your answer in grams, rounded to the nearest gram.

OpenStudy (jadedry):

This is the balanced equation. \[C _{3}H _{8} +5O_{2} -> 3CO _{2} + 4H _{2}O\] The molar mass of \[C _{3}H _{8}\] is: 3*12 + 8*1 = 44g Therefore you have 132/44 = 3 moles of propane. As you are given enough oxygen to fully burn the propane, the equation becomes: \[3(C _{3}H _{8}) + 15(O _{2} )= 9(CO _{2}) + 12(H _{2}O)\] Therefore, you will have 12 moles of water. The weight of a mole of water = 16 + 1*2 = 18g To find the total weight of water produced, 12 moles * 18g = 216 grams

OpenStudy (anonymous):

Thank you so much for helping me and explaining it. I didn't understand what I was doing but again thanks c:

OpenStudy (jadedry):

No problem! c: